A pH 4.01 buffer is to be prepared by combining 0.125 M acetic acid with 0.150 M sodium acetate.  How many mL of each substance is needed to prepare 250.0 mL of this buffer?
Answer #1

Let volume of acetic acid needed = x mL

then, volume of sodium acetate needed = (total volume) - (volume of acetic acid)

volume of sodium acetate needed = 250 mL - x mL

volume of sodium acetate needed = (250 - x) mL

moles acetic acid needed = (concentration acetic acid) * (volume of acetic acid needed)

moles acetic acid needed = (0.125 M) * (x mL)

moles acetic acid needed = 0.125x mmol

moles sodium acetate needed = (concentration sodium acetate) * (volume of sodium acetate needed)

moles sodium acetate needed = (0.150 M) * (250 - x) mL

moles sodium acetate needed = (37.5 - 0.150x) mmol

Ka acetic acid = 1.8 x 10-5

pKa = -log(Ka)

pKa = -log(1.8 x 10-5)

pKa = 4.74

According to Henderson - Hasselbalch equation,

pH = pKa + log([conjugate base] / [weak acid])

pH = pKa + log(moles sodium acetate needed / moles acetic acid needed)

Substituting the values,

4.01 = 4.74 + log((37.5 - 0.150x) / 0.125x)

log((37.5 - 0.150x) / 0.125x) = 4.01 - 4.74

log((37.5 - 0.150x) / 0.125x) = -0.73

(37.5 - 0.150x) / 0.125x = 10-0.73

(37.5 - 0.150x) / 0.125x = 0.184

37.5 - 0.150x = (0.184) * (0.125x)

37.5 - 0.150x = 0.023x

37.5 = 0.150x + 0.023x

37.5 = 0.173x

x = (37.5) / (0.173)

x = 216.7

volume of acetic acid needed = x mL

volume of acetic acid needed = 216.7 mL

volume of sodium acetate needed = (250 - x) mL

volume of sodium acetate needed = (250 - 216.7) mL

volume of sodium acetate needed = 33.3 mL